trying to help my 8th grade granddaughgter but dont know what to do
need width of rectangle with area 24 inches to 2 power if length is 1 inch

Answer by mananth(16946)

if i understand rightly what is written
area = 24^2 sq.inches
length = 1 inches
...
Area = L*W
24^2 = 1*W
w= 24 ^2 inches
= 576 inches
...
m.ananth@hotmail.ca